From nch@roe.ac.uk Fri May 15 17:01:49 2009 Date: Fri, 15 May 2009 16:45:06 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 15/05/09 Minutes of WFAU VDFS Science Archive meeting: 15th May 2009 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, ETWS, NJC, MSH, MAR, RGM, RPB, RSC Apologies: JPE, JMS, AL, PMW DoNM: 10am, Friday May 15th 2009 in the Vista Hut Actions discharged this week: ----------------------------- None this week. Actions partly discharged but continuing: ----------------------------------------- ACTION: ETWS to parse that latest keyword info into the VSA metadata schema files. Continues; ETWS noted that communications with CASU indicate that the VIRCam data FITS keys are yet to be finalised, but that the procedure for parsing a standard set into an SQL schema is set up and can be tested further in the meantime. Actions carried forward from 8/05/09 meeting: ---------------------------------------------- None this week Specific points and new actions: -------------------------------- Project management: ACTION: NCH to make polite enquiries concerning grant renewal status to STFC attendees at the UKIRT Board meeting next week WFCAM & VISTA updates: WFCAM continues to pump out lots of data. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: ETWS noted that WFAU's login at CASU is now back up and running, and also that we're anticipating the first tranche of 09A data next week (thanks to MJI for the communications). WSA Operations: RPB noted: "Non surveys continue. Trying to release U07BH60 again, as that caused us big problems last time trying to calibrate ZeroPoint. It's not the fastest non survey as it's a very crowded field, but it's still plodding along. Attempting to split up the monolithic BestDR7 into more manageable chunks. NCH's attempts to run the supplied scripts ground to a halt, so we're trying again with DTS. Fingers crossed!" Hardware: RPB noted: "Another disk failure on djedefre, on the same disk slot. Replaced the disk, though the new one wasn't recognised by the system. Reseated all cables associated with that disk and it came up OK. RAID array is now rebuilt and happy, though suggest we keep an eye on that one and request new RAID card from eclipse should it fail again." Software: ETWS noted finishing the upgrade of LGR's SDSSPhotometry code and started working on the UKIDSSPhotometry. NCH noted that some feedback on the tools has been coming in from a few users, so it's good to know they will be useful. RSC asked about the SW installation on 64-bit workhorse server menkaure, and ETWS confirmed that this was now complete and ready for testing. (MSH noted that he has yet to upgrade the physical memory of that machine). RSC reported: "Working with NCH on modifications to CU7 frame set association to handle the new second-epoch requirements of the GPS survey for the DR5 release. Also took the opportunity to do further investigations on the frame set driven design of CU7 to see if it can be made more robust against failure during large runs. Prior to working on CU7 I had got stuck into the code changes for the dual VSA & WSA curation design, which have not been committed to SVN yet." NJC reported: "This week I have been finishing off the paper on archiving multi-epoch data. I have sent a second draft around to the WFAU team and expect to submit on Monday." Survey Data NCH noted that we're now ready to source merge the GPS for DR5 thanks to the software mods done by RSC; NCH will also find out the status of the UDS DR5 stacking next week when he's down south for the UKIRT Board meeting in Nottingham. Non-survey Data Release: No new releases this week. Astrogrid deployment & Data Analysis services: MSH noted that he has now got the VOTable to KML application working on the VODesktop, and MAR has been experimenting with colour images, all for GoogleSky overlays. Miscellaneous: Nothing else this week. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================